OpenStudy (anonymous):

solve for 1 ------- a+bi dont you times the top with a-bi

OpenStudy (anonymous):

You times the top and bottom by a - bi

OpenStudy (anonymous):

1(a - bi) ------------ (a + bi)(a - bi)

OpenStudy (anonymous):

a - bi ---------- (a^2 + b^2) Do you understand the denominator..... The -abi and the +abi cancel out. The (bi)(-bi) = -b^2i^2 but i^2 = -1 so you have -b^2(-1) = b^2